# Domain-Invariant Representation Learning of Bird Sounds

### v3

@@ Abstract
Passive acoustic monitoring (PAM) is crucial for bioacoustic research, enabling non-invasive species tracking and biodiversity monitoring. Citizen science platforms like Xeno-Canto provide large annotated datasets from focal recordings, where the target species is intentionally recorded. However, PAM requires monitoring in passive soundscapes, creating a domain shift between focal and passive recordings, which challenges deep learning models trained on focal recordings. To address this, we leverage supervised contrastive learning to improve domain generalization in bird sound classification, enforcing domain invariance across same-class examples from different domains. We also propose ProtoCLR (Prototypical Contrastive Learning of Representations), which reduces the computational complexity of the SupCon loss by comparing examples to class prototypes instead of pairwise comparisons. Additionally, we present a new few-shot classification benchmark based on BirdSet, a large-scale bird sound dataset, and demonstrate the effectiveness of our approach in achieving strong transfer performance. [^1] Models and code: https://github.com/ilyassmoummad/ProtoCLR
@@ I Introduction
We propose a new few-shot classification benchmark based on the BirdSet [12] dataset to evaluate the generalization capabilities of models trained on XC’s focal recordings and tested on diverse soundscape datasets. This benchmark is designed to assess how well models can generalize across domains in challenging few-shot scenarios. We validate our proposed loss ProtoCLR on this benchmark, demonstrating its effectiveness in improving DG in bird sound classification.
@@
  • We establish a large-scale few-shot benchmark for bird sound classification using BirdSet datasets, evaluating model generalization from focal to soundscape recordings.
@@ IV Experiments
Split Dataset # of recordings # of 5s examples # of classes
Train XCL 528,434 1,401,478 9,736
XCM 89,798 189,880 411
Val POW 14,911 51,394 48
Test PER 16,802 65,479 132
NES 16,117 58,251 89
UHH 3,626 12,980 27
HSN 5,460 17,940 21
NBP 24,327 76,446 51
SSW 28,403 92,514 81
SNE 19,390 64,978 56
TABLE I: BirdSet Classification Benchmark.
Val Test
Model POW PER NES UHH HSN NBP SSW SNE Mean
Random Guessing 2.08 0.75 1.12 3.70 4.76 1.96 1.23 1.78 2.17
@@
BirdAVES 7.55$\pm$0.8 8.38$\pm$0.5 8.83$\pm$1.1 24.1$\pm$4.4 13.5$\pm$3.0 10.2$\pm$1.3 6.64$\pm$0.9 7.19$\pm$1.1 10.8
BioLingual 39.6$\pm$4.6 33.8$\pm$1.5 41.2$\pm$4.0 59.2$\pm$4.6 50.3$\pm$9.8 44.8$\pm$3.1 39.9$\pm$4.0 41.6$\pm$0.4 43.8
Perch 39.4$\pm$3.0 41.7$\pm$0.8 45.0$\pm$3.6 59.3$\pm$8.1 46.5$\pm$7.3 48.1$\pm$2.9 40.5$\pm$4.7 40.7$\pm$8.6 45.1
XCM Pre-training
CE 43.6$\pm$3.5 41.8$\pm$1.4 45.2$\pm$2.7 67.2$\pm$6.5 58.8$\pm$4.9 34.4$\pm$2.9 37.5$\pm$3.2 41.3$\pm$3.6 46.2
SimCLR 15.5$\pm$3.2 12.4$\pm$0.9 14.9$\pm$1.4 27.2$\pm$4.7 22.0$\pm$3.4 15.1$\pm$1.3 12.0$\pm$1.5 14.3$\pm$1.8 16.6
SupCon 69.3$\pm$4.4 75.3$\pm$2.1 72.8$\pm$4.0 81.2$\pm$8.4 74.8$\pm$7.1 58.4$\pm$3.9 62.5$\pm$4.5 62.6$\pm$4.1 69.6
ProtoCLR 72.3$\pm$4.8 79.0$\pm$2.2 79.1$\pm$2.9 83.4$\pm$6.2 82.2$\pm$6.2 65.2$\pm$3.5 69.6$\pm$3.5 69.6$\pm$5.9 75.0
XCL Pre-training
CE 32.3$\pm$2.9 27.6$\pm$1.9 34.1$\pm$2.5 54.5$\pm$7.4 44.8$\pm$4.4 31.6$\pm$1.8 27.7$\pm$2.1 30.4$\pm$2.6 35.3
SimCLR 15.0$\pm$1.9 12.8$\pm$0.8 15.3$\pm$1.6 27.6$\pm$5.7 19.6$\pm$3.6 15.4$\pm$1.3 12.7$\pm$1.9 15.1$\pm$2.3 16.7
SupCon 35.4$\pm$5.5 36.1$\pm$2.1 37.1$\pm$3.5 58.7$\pm$4.2 45.1$\pm$5.2 60.7$\pm$4.1 40.0$\pm$3.0 44.4$\pm$4.6 44.7
ProtoCLR 30.7$\pm$3.2 28.9$\pm$1.5 32.3$\pm$2.6 51.5$\pm$4.5 39.4$\pm$4.4 50.6$\pm$3.4 33.5$\pm$3.5 38.6$\pm$4.1 38.2
@@
BirdAVES 15.5$\pm$0.8 18.8$\pm$0.7 18.5$\pm$1.0 37.0$\pm$7.0 23.8$\pm$2.2 18.5$\pm$1.2 13.8$\pm$1.1 15.8$\pm$1.0 20.2
BioLingual 65.8$\pm$0.9 58.8$\pm$0.5 66.0$\pm$0.7 77.5$\pm$1.5 72.0$\pm$2.9 70.7$\pm$0.5 65.4$\pm$0.8 64.7$\pm$1.7 67.6
Perch 67.3$\pm$1.2 68.7$\pm$0.7 71.8$\pm$0.6 82.0$\pm$1.8 78.7$\pm$1.6 79.0$\pm$0.5 69.8$\pm$1.0 72.1$\pm$1.0 73.7
XCM Pre-training
CE 79.8$\pm$0.9 80.7$\pm$0.6 82.6$\pm$0.8 91.2$\pm$1.4 88.6$\pm$1.0 63.3$\pm$1.0 74.9$\pm$1.1 78.1$\pm$1.4 79.9
SimCLR 28.8$\pm$1.4 27.3$\pm$0.7 30.0$\pm$1.2 47.1$\pm$3.5 38.3$\pm$1.7 28.3$\pm$1.1 25.0$\pm$1.2 28.6$\pm$2.1 31.7
SupCon 83.8$\pm$0.5 87.7$\pm$0.2 86.4$\pm$0.3 90.2$\pm$0.8 88.6$\pm$0.7 77.1$\pm$0.4 80.8$\pm$0.7 80.9$\pm$1.0 84.4
ProtoCLR 87.5$\pm$0.4 89.9$\pm$0.2 89.2$\pm$0.3 91.8$\pm$0.8 92.2$\pm$0.5 81.0$\pm$0.6 85.0$\pm$0.6 85.8$\pm$0.8 87.8
XCL Pre-training
CE 65.8$\pm$1.3 65.3$\pm$0.7 71.3$\pm$1.0 86.6$\pm$1.8 79.1$\pm$1.7 64.6$\pm$1.3 64.1$\pm$1.1 67.5$\pm$1.6 70.5
SimCLR 30.0$\pm$1.4 28.3$\pm$0.9 31.0$\pm$1.0 49.5$\pm$3.4 38.0$\pm$1.7 30.2$\pm$1.0 26.7$\pm$1.4 29.2$\pm$1.8 32.8
SupCon 58.3$\pm$1.2 57.8$\pm$0.7 60.6$\pm$0.8 76.4$\pm$1.7 69.5$\pm$2.2 77.2$\pm$0.5 60.2$\pm$1.0 62.9$\pm$1.0 65.3
ProtoCLR 52.3$\pm$1.7 51.2$\pm$0.7 54.2$\pm$1.3 73.1$\pm$1.1 63.0$\pm$2.1 71.8$\pm$0.9 53.0$\pm$1.2 55.9$\pm$1.4 59.3
TABLE II: Top-1 accuracy for one-shot and five-shot classification. All reported results are the average of ten runs. For our models, the results are further averaged over three checkpoints. The highest accuracies are highlighted in bold, and the second highest are underlined.
@@ IV-A Few-Shot Classification Benchmark
The BirdSet benchmark [12] comprises two tasks: multi-label classification, where each audio recording is segmented into 5-second intervals to detect the presence of one or more species (or none), and multi-class classification, where individual bird events are detected using peak detection and bambird [23], a tool for identifying bird events in audio recordings, with each event containing a single species. In line with the MetaAudio framework, we focus on the multi-class classification task to define a few-shot evaluation for assessing the generalization capabilities of pre-trained models [24]. BirdSet offers two training sets: XCL (XC Large) with focal recordings across nearly 10,000 species, and XCM (XC Medium), a specialized subset of XCL with recordings from 411 species represented in the test datasets. The validation and test datasets contain soundscape recordings. The benchmark is detailed in Table I.
@@ IV-B Reference Systems
To compare our domain-invariant pre-training approach (SupCon and ProtoCLR), we train reference systems using cross-entropy (CE) loss as a supervised baseline and SimCLR as a self-supervised contrastive baseline. Additionally, we evaluate large-scale, state-of-the-art models in bioacoustics: BirdAVES [27], a 24-layer transformer with 1024 units, following the same architecture as HuBERT-large [28]. It is trained on bird sounds from datasets such as AudioSet, VGGSound, XC, and iNaturalist; the encoder of BioLingual [29], an HT-SAT transformer pre-trained on AudioSet and fine-tuned using contrastive language-audio training to align animal sounds with text captions describing the class across a large collection of data including XC, iNaturalist, Animal Sound Archive, …etc; and the encoder of Perch [16], an EfficientNet-B1 [30] trained on XC for species classification, as well as taxonomic ranks genus, family, and order.
@@ IV-C Pre-training Details
We train all models with CvT-13 [31], an efficient transformer architecture, on XCM and XCL datasets for 100 epochs using the AdamW optimizer with a batch size of 256, with a weight decay of $1\times 10^{-6}$. Following Moummad et al.[6], we apply the augmentations found to be effective for bird sound representations: circular time shift[32], SpecAugment [33], and spectrogram mixing [34]. These models are trained with a projector of dimension 128. For the CE loss, we only apply circular time shift and SpecAugment as augmentations, excluding Spectrogram Mixing, as it prevented the model from converging. The learning rate for CE and ProtoCLR is set to $5\times 10^{-4}$, while for SupCon and SimCLR, we use a learning rate of $1\times 10^{-4}$. We tune hyperparameters by monitoring k-NN accuracy on the POW dataset.
@@ IV-D Results and Discussion
Table II presents the performance of different models on one-shot and five-shot bird sound classification tasks. ProtoCLR pre-trained on XCM consistently outperforms others in both tasks, with SupCon close behind. Additionally, ProtoCLR is more computationally efficient; for one training epoch with a batch size of 256, SupCon computes 80.4B MACs, while ProtoCLR computes only 28.3B. Notably, ProtoCLR significantly outperforms CE on average (75.0 vs. 46.2).
For self-supervised methods, SimCLR outperforms BirdAVES in both one-shot and five-shot learning, suggesting that invariant learning is more effective than self-prediction methods for few-shot scenarios, consistent with findings in computer vision [35]. However, among our trained models, SimCLR performs the worst on XCM, likely due to domain shift. Incorporating unsupervised domain generalization techniques [10] could enhance SimCLR’s generalization ability.
Pre-training on the larger XCL dataset negatively impacts the performance of ProtoCLR, SupCon, and CE models. This may be due to XCL’s broader class diversity (9,736 classes) compared to XCM (411 classes) that only contains the target classes, which leads to a loss of discriminative capacity on the target classes. XCL also has a higher coefficient of variation (1.43 vs 0.43) and Gini coefficient (0.62 vs 0.24), indicating greater class imbalance, which ProtoCLR may be more sensitive to than SupCon.
When pretrained on XCL, SupCon and ProtoCLR outperform CE in one-shot learning but not in five-shot learning, where SupCon surpasses ProtoCLR. Additionally, SupCon performs comparably to BioLingual on both one-shot and five-shot tasks. Perch slightly outperforms CE on XCL, suggesting that incorporating taxonomic ranks as auxiliary tasks could be a promising direction for future research.
These results indicate that ProtoCLR and SupCon achieves strong few-shot performance especially in one-shot learning scenarios.
@@ V Conclusion
In this work, we addressed the challenge of domain generalization for bird sound classification in few-shot scenarios, focusing on the domain shift from focal to soundscape recordings. We proposed a new few-shot benchmark derived from the BirdSet dataset to evaluate generalization capabilities of models trained on focal recordings and tested on soundscape recordings. Additionally, we introduced ProtoCLR, an alternative to SupCon inspired by prototypical networks, with reduced computational complexity. Few-shot evaluation under two pre-training scenarios—one with recordings limited to test classes and another with a larger, more diverse set—showed that pre-training on the set containing only test classes leads to better transfer performance.
Future work could explore techniques to scale to larger class sets without compromising performance, as well as investigate domain adaptation methods, particularly source-free adaptation, which are suited for few-shot bioacoustics when source data is unavailable or too costly to retrain on.

### v4

@@ Abstract
Passive acoustic monitoring (PAM) is crucial for bioacoustic research, enabling non-invasive species tracking and biodiversity monitoring. Citizen science platforms like Xeno-Canto provide large annotated datasets from focal recordings, where the target species is intentionally recorded. However, PAM requires monitoring in passive soundscapes, creating a domain shift between focal and passive recordings, which challenges deep learning models trained on focal recordings. To address this, we leverage supervised contrastive learning to improve domain generalization in bird sound classification, enforcing domain invariance across same-class examples from different domains. We also propose ProtoCLR (Prototypical Contrastive Learning of Representations), which reduces the computational complexity of the SupCon loss by comparing examples to class prototypes instead of pairwise comparisons. Additionally, we present a new few-shot classification evaluation based on BIRB, a large-scale bird sound benchmark to evaluate bioacoustic pre-trained models. [^1] Models and code: https://github.com/ilyassmoummad/ProtoCLR
@@ I Introduction
We propose a new few-shot classification evaluation based on the BIRB [12] benchmark to evaluate the generalization capabilities of models trained on XC’s focal recordings and tested on diverse soundscape datasets. This benchmark is designed to assess how well models can generalize across domains in challenging few-shot scenarios. We validate our proposed loss ProtoCLR on this benchmark, demonstrating its effectiveness in improving DG in bird sound classification.
@@
  • We establish a large-scale few-shot benchmark for bird sound classification using BIRB datasets, evaluating model generalization from focal to soundscape recordings.
@@ IV Experiments
Split Dataset # of recordings # of classes
Train XC 684,744 10,127
Val POW 16,052 48
Test PER 14,798 132
NES 6,952 89
UHH 59,583 27
HSN 10,296 19
SSW 50,760 96
SNE 20,147 56
TABLE I: Overview of BIRB datasets.
Model PER NES UHH HSN SSW SNE Mean
Random Guessing 0.75 1.12 3.70 5.26 1.04 1.78 2.22
@@
BirdAVES-biox-base 7.41$\pm$1.0 26.4$\pm$2.3 13.2$\pm$3.1 9.84$\pm$3.5 8.74$\pm$0.6 14.1$\pm$3.1 13.2
BirdAVES-bioxn-large 7.59$\pm$0.8 27.2$\pm$3.6 13.7$\pm$2.9 12.5$\pm$3.6 10.0$\pm$1.4 14.5$\pm$3.2 14.2
BioLingual 6.21$\pm$1.1 37.5$\pm$2.9 17.8$\pm$3.5 17.6$\pm$5.1 22.5$\pm$4.0 26.4$\pm$3.4 21.3
Perch 9.10$\pm$5.3 42.4$\pm$4.9 19.8$\pm$5.0 26.7$\pm$9.8 22.3$\pm$3.3 29.1$\pm$5.9 24.9
Ours
CE 10.6$\pm$1.2 35.0$\pm$4.1 18.0$\pm$3.3 19.5$\pm$5.3 13.3$\pm$1.5 22.9$\pm$5.1 19.8
SimCLR 8.61$\pm$1.1 30.6$\pm$3.4 13.8$\pm$2.7 15.3$\pm$5.0 9.11$\pm$0.9 17.1$\pm$2.7 15.7
SupCon 7.87$\pm$0.9 34.5$\pm$2.5 15.6$\pm$3.4 18.5$\pm$4.9 12.1$\pm$1.5 19.6$\pm$3.7 18.0
ProtoCLR 8.30$\pm$1.1 36.5$\pm$3.4 15.2$\pm$3.0 19.8$\pm$5.6 11.7$\pm$1.6 20.3$\pm$3.3 18.6
@@
BirdAVES-biox-base 11.6$\pm$0.8 39.7$\pm$1.8 22.5$\pm$2.4 22.1$\pm$3.3 16.1$\pm$1.7 28.3$\pm$2.3 23.3
BirdAVES-bioxn-large 15.0$\pm$0.9 42.6$\pm$2.7 23.7$\pm$3.8 28.4$\pm$2.4 18.3$\pm$2.3 27.3$\pm$2.3 25.8
BioLingual 13.6$\pm$1.3 65.2$\pm$1.4 31.0$\pm$2.9 34.3$\pm$3.5 43.9$\pm$0.9 49.9$\pm$2.3 39.6
Perch 21.2$\pm$1.2 71.7$\pm$1.5 39.5$\pm$3.0 52.5$\pm$5.9 48.0$\pm$1.9 59.7$\pm$1.8 48.7
Ours
CE 22.0$\pm$1.0 64.6$\pm$3.1 34.9$\pm$2.4 46.6$\pm$4.5 30.6$\pm$1.1 46.6$\pm$2.1 40.8
SimCLR 17.4$\pm$0.9 52.1$\pm$2.3 23.4$\pm$2.4 27.2$\pm$3.6 19.7$\pm$1.4 30.5$\pm$2.2 28.4
SupCon 16.1$\pm$1.0 60.3$\pm$2.0 29.7$\pm$2.5 35.7$\pm$5.6 25.9$\pm$1.2 39.4$\pm$2.7 34.5
ProtoCLR 16.5$\pm$1.1 61.3$\pm$2.3 30.0$\pm$2.4 38.4$\pm$4.4 26.0$\pm$1.3 39.5$\pm$2.3 35.3
TABLE II: Top-1 accuracy for one-shot and five-shot classification. All reported results are the average of ten runs. For our models, the results are further averaged over three checkpoints.
@@ IV-A Few-Shot Classification Benchmark
The original task of BIRB [12] is information retrieval in bioacoustics, focused on retrieving bird vocalizations from passively recorded datasets using focal recordings for training. Downstream datasets may contain long audio recordings where events of interest are sparse. Individual bird events are detected using peak detection [23], which identifies frames with high energy to used for the task. We build on our previous work [6], which shows that using a pretrained classifier on AudioSet [24] is a good proxy for selecting windows with the highest bird class activation.
Following the MetaAudio framework, we focus on a multi-class classification task to define a few-shot evaluation for assessing the generalization capabilities of pre-trained models [25]. BIRB provides a large training set, XC (Xeno-Canto), consisting of focal recordings from nearly 10,000 species, while the validation and test datasets contain soundscape recordings. The benchmark details are provided in Table I.
@@ IV-B Reference Systems
To compare supervised contrastive approaches SupCon and ProtoCLR for addressing domain generalization, we train reference systems using cross-entropy (CE) loss as a supervised baseline and SimCLR as a self-supervised contrastive baseline. Additionally, we evaluate large-scale, state-of-the-art models in bioacoustics: BirdAVES [28], a transformer model based on the speech model HuBERT, trained in a self-supervised manner based on the AVES framework (explained in Section LABEL:motiv). We evaluate two versions of BirdAVES, biox-base, a 12-layer transformer with 768 units, trained on Xeno-Canto [xenocanto] and animal sounds from AudioSet [24] and VGGSound [vggsound], and bioxn-large, a 24-layer transformer with 1024 units, trained on the same datasets plus iNaturalist  [^2] https://github.com/gvanhorn38/iNatSounds. BirdAVES outputs one embedding vector per time frame, which we average across the time dimension to produce a single vector representing the entire audio clip; the encoder of BioLingual [29], an HT-SAT transformer pre-trained on AudioSet and fine-tuned using contrastive language-audio training to align animal sounds with text captions describing the class across a large collection of data including Xeno-Canto, iNaturalist, Animal Sound Archive, …etc; and the encoder of Perch [12], an EfficientNet-B1 [30] trained on Xeno-Canto for species classification, as well as taxonomic ranks genus, family, and order. We evaluate only models trained specifically on bird sounds, as general-purpose audio models have been shown to perform poorly in this domain [25]. Similarly, Hamer et al.[12] found that general-purpose models can, in some cases, perform worse than simply averaging mel spectrogram features without any learning.
@@ IV-C Pre-training Details
We follow the same preprocessing as Moummad et al.[6]. We train all models with CvT-13 [31], an efficient 2D transformer architecture with 20M parameters, on XC for 100 epochs using the AdamW optimizer with a batch size of 256, with a weight decay of $1\times 10^{-6}$. Following Moummad et al.[6], we apply the augmentations found to be effective for bird sound representations: circular time shift [32], SpecAugment [33], and spectrogram mixing [34]. These models are trained with a projector of dimension 128. For the CE loss, we only apply circular time shift and SpecAugment as augmentations, excluding Spectrogram Mixing, as it prevented the model from converging. The learning rate for CE and ProtoCLR is set to $5\times 10^{-4}$, while for SupCon and SimCLR, we use a learning rate of $1\times 10^{-4}$. We tune hyperparameters by monitoring k-NN accuracy on the POW dataset, which is split randomly into a training and a validation subset.
@@ IV-D Results and Discussion
Table II presents the performance of different models on one-shot and five-shot bird sound classification tasks. ProtoCLR slightly outperforms SupCon both in one-shot and five-shot classification. Additionally, ProtoCLR is more computationally efficient; for one training epoch with a batch size of 256, SupCon computes 80.4B MACs, while ProtoCLR computes only 28.3B. On the other hand, CE outperforms ProtoCLR, when trained on 100 epochs. This could be due to the fact that contrastive methods require longer training iterations for optimal convergence than cross-entropy [5], which should be investigated.
For self-supervised methods, SimCLR outperforms BirdAVES-biox-base and BirdAVES-bioxn-large in both one-shot and five-shot learning, suggesting that invariant learning is more effective than self-prediction methods for few-shot scenarios, consistent with findings in computer vision [35].
On average, BioLingual outperforms CE in one-shot learning but underperforms in five-shot learning. However, when examining individual datasets, BioLingual performs worse on PER, NES, UHH, and HSN, yet significantly better on SSW and SNE. Interestingly, certain models appear to specialize in specific species, suggesting a promising research direction: distilling these specialized models into a single general model that can leverage the strengths of each.
Perch significantly outperforms all models in both one-shot and five-shot classification across all datasets, except for PER, where it performs slightly worse than CE. This suggests that incorporating taxonomic ranks as auxiliary tasks may enhance the discriminative capabilities of the feature extractor.
@@ V Conclusion
In this work, we addressed the challenge of domain generalization for bird sound classification in few-shot scenarios, with a focus on the domain shift from focal to soundscape recordings. We proposed a new few-shot evaluation, derived from the BIRB datasets, to evaluate the generalization capabilities of models trained on focal recordings and tested on soundscapes. Additionally, we introduced ProtoCLR, a computationally efficient alternative to SupCon, inspired by prototypical networks. ProtoCLR maintains performance levels comparable to SupCon, and in some cases, slightly outperforms it.
Future work will explore scaling training to longer epochs to maximize the potential of contrastive methods, as well as integrating taxonomic ranks into the learning process. Another promising direction is to investigate domain adaptation techniques, particularly source-free adaptation, which is well-suited for few-shot bioacoustics when access to source data is limited or retraining is too costly.